# Delving Deeper Into Astromorphic Transformers

### 18 Dec 2023 v1

@@ Results
Contribution II: Prior work (Kozachkov, Kastanenka, and Krotov 2023) at modelling pre-synaptic plasticity has not considered the intrinsic non-linear $Ca^{2+}$ dynamics of the astrocytes in response to increasing firing rate of presynaptic neurons. Due to the indirect feedback (eSP), the rate of increase in the release rate of neurotransmitters gradually reduces in response to an increase in presynaptic neuron firing over time. Hence, a non-linear relationship exists between the calcium concentration and the frequency of spikes received from the presynaptic neuron. We substantiate this observation by simulating the AN model (Wade et al. 2012b, 2011). The associated non-linear variation is depicted in Fig. 3 which leads us to modify the presynaptic plasticity parameter $g$ as,
@@
Realization of Self-Attention
@@
Let us define the weight modulating the connection between the astrocyte and the output layer as $P$. According to studies regarding Tumor Necrosis Factor - $\alpha$ ($TNF-\alpha$), astrocytes possess the ability to boost synaptic weights in reaction to low neural activity and diminish weights in reaction to high neural activity (Kozachkov, Kastanenka, and Krotov 2023) by means of calcium response (neural activity is proportional to released neurotransmitters). Consequently, it is possible to infer an inverse correlation between $P$ and the calcium response ($C$). Simplifying, this relation is given by,
@@
Table 1: Performance comparison of our proposed astromorphic attention-based transformer architecture against other benchmarks with varying degrees of bioplausibility on the IMDB and CIFAR10 datasets.
@@
Performance Evaluation
Our astromorphic transformer is trained and tested on two different tasks. These include sentiment classification on the IMDB dataset (Maas et al. 2011) and image classification on the CIFAR10 dataset (Krizhevsky and Hinton 2009). Methods section discusses the hardware specifications, datasets and model architectures while Appendix C in the supplementary material lists all the associated hyper-parameters.
To date, there has not been any language or vision model that incorporates transformer architecture in a manner that is bioplausible and derived from neuron-astrocyte interactions. We report an accuracy of $88.7\pm 0.2\%$ for the sentiment classification task on the IMDB dataset, utilizing the non-linearity and relative positional information in the model. We also achieve $97.0\pm 0.2\%$ accuracy using an astromorphic vision transformer on the CIFAR10 dataset.
Table 1 compares performance of our astromorphic transformer against neuromorphic as well as other non-bioplausible models on the IMDB and CIFAR10 datasets. Our model outperforms RNNs, LSTMs, and GRUs on the IMDB dataset and other bioplausible neuromorphic architectures. However, we did not compare with existing pretrained transformers (BERT, GPT, etc.) because these models are pretrained using non-bioplausible self-attention and have customized pretrained learnable embeddings for natural language processing, whereas we use fixed GloVe embeddings (Pennington, Socher, and Manning 2014) that are simple enough to implement in our astromorphic attention-based transformer architecture. We also test the competitiveness of astromorphic attention in vision transformer architecture. As shown in Table 1, our astromorphic vision transformer model competes with artificial and outperforms bioplausible models on the CIFAR10 dataset.
Ablation Studies
Learning speed:
@@
Final accuracy:
We implement the theoretical model proposed by (Kozachkov, Kastanenka, and Krotov 2023) to see how it compares without the bioplausibilities proposed herein and find an accuracy of $88.0\pm 0.2\%$ and $87.4\pm 0.1\%$ on the IMDB and CIFAR10 datasets, respectively. This implementation does not incorporate the non-linearity and relative positional information that are integral to our model, which draws inspiration from the dynamics between neurons and astrocytes, thereby underscoring the strong need to consider such bioplausibilities in modeling astromorphic transformers. Our astromorphic transformer also competes with an iso-architecture standard non-bioplausible transformer with softmax attention (see “Transformer” entry in Table 1).
Discussion
@@
Methods
Hardware Specifications
Dataset Details
The IMDB dataset, created by Stanford University, contains $50,000$ movie reviews (each 250-500 words long) labeled as positive or negative. The data is presented in a tab-delimited format, containing fields for movie title, review text, and sentiment label. The CIFAR-10 dataset is a collection of computer vision data, comprising $60,000$ color images with dimensions of $32\times 32$ pixels. These images are categorized into 10 distinct classes, with each class containing 6,000 images. The dataset exhibits a relatively diminutive size, comprising low-resolution images and notable diversity within each class.
Network Architecture
Meanwhile, for image classification, a convolutional and self-attention based neural network (CoAtNet (Dai et al. 2021)) was utilized to implement our proposed innovations. CoAtNet transformer architecture can best utilize the effectiveness of relative positional information when trained on vision task focused datasets. Furthermore, CoAtNet leverages the capabilities of Convolutional Neural Networks (CNNs) and transformers to enhance its performance. Our architecture implements a $3\times 3$ convolution before creating patches from the images. We use a patch size of $2$ on the image inputs and employ an 8-layered encoder stack with 8-headed self-attention. In addition to the original CoAtNet architecture, we add additional non-linearity and average pooling after encoder layers.
@@ Methods
Appendix A: Transformer Illustration
Transformer employs only encoder architecture for performing sentiment classification and vision tasks. As illustrated in Fig. 5, the input embeddings combined with the positional embeddings enter the encoder layer as input tokens. Then key, query and value are generated from the input tokens and after performing self attention on them, a skip connection is added from the inputs and the overall result is layer normalized and fed into a feed forward network. Another add & normalize operation is performed before feeding them into a linear layer and softmax function to obtain the output probabilities. Sometimes, there might be more than one encoder layer. For different tasks, like machine translation and text generation, encoder-decoder or only decoder type architecture of transformer can be used. However, independent of the type of architecture needed, self-attention is a fundamental mechanism defining transformer operation. The “Multi-Head Attention” Block has self-attention modules parallelly running on different heads obtained from the input embeddings.
Figure 5: Schematic representation of the Transformer Encoder architecture, illustrating the sequential flow of operations. $N$ denotes the number of encoder layers in the network.
Appendix B: Memory Efficient Linearized Attention
Linearized attention involves utilizing feature maps with lower complexity levels, thereby minimizing memory and computational resources. As discussed before, the introduction of non-linearity in biological neurons is facilitated by an activation function ($elu(X)$) in the hidden layer neurons. The utilization of the positive similarity feature map, denoted by $\phi(.)$, characterizes the implementation of the linearized attention operation as described in the work by (Katharopoulos et al. 2020). The utilization of linearized feature attention results in a reduction of computational complexity associated with the computation of the attention score. The advantage of utilizing feature map based linearized attention over the conventional distribution function (i.e. softmax) based self-attention is demonstrated in Fig. 6.
@@
Table 2: Hyper-parameter list for the IMDB dataset.
Figure 6: Figure showing the comparison of order of complexity between usual and linearized self-attention. Distribution function can be mapped to softmax function used in usual transformers.
Appendix C: Hyper-parameter Details
The hyper-parameters for the encoder-type astromorphic transformer for the IMDB dataset are listed in Table 2 in the supplementary material. The hyper-parameters for the encoder-type convolution and astromorphic self-attention based transformer for the CIFAR10 dataset are listed in Table 3.
@@
Table 3: Hyper-parameter list for the CIFAR10 dataset.

### 09 Jan 2024 v2

@@ III-D2 Presynaptic Plasticity
Contribution II: Prior work [18] at modelling pre-synaptic plasticity has not considered the intrinsic non-linear $Ca^{2+}$ dynamics of the astrocytes in response to increasing firing rate of presynaptic neurons. Due to the indirect feedback (eSP), the rate of increase in the release rate of neurotransmitters gradually reduces in response to an increase in presynaptic neuron firing over time. Hence, a non-linear relationship exists between the calcium concentration and the frequency of spikes received from the presynaptic neuron. We substantiate this observation by simulating the AN model [40, 37]. The associated non-linear variation is depicted in Fig. 3 which leads us to modify the presynaptic plasticity parameter $g$ as,
@@
III-E Realization of Self-Attention
@@
Let us define the weight modulating the connection between the astrocyte and the output layer as $P$. According to studies regarding Tumor Necrosis Factor - $\alpha$ ($TNF-\alpha$), astrocytes possess the ability to boost synaptic weights in reaction to low neural activity and diminish weights in reaction to high neural activity [18] by means of calcium response (neural activity is proportional to released neurotransmitters). Consequently, it is possible to infer an inverse correlation between $P$ and the calcium response ($C$). Simplifying, this relation is given by,
@@
TABLE I: Performance comparison of our proposed astromorphic attention-based transformer architecture against other benchmarks with varying degrees of bioplausibility on the IMDB and CIFAR10 datasets.
@@
IV Results
Our astromorphic transformer is trained and tested on two different tasks. These include sentiment classification on the IMDB dataset [58] and image classification on the CIFAR10 dataset [59]. Appendix C in the supplementary material lists all the associated hyper-parameters. To date, there has not been any language or vision model that incorporates transformer architecture in a manner that is bioplausible and derived from neuron-astrocyte interactions. We report an accuracy of $88.7\pm 0.2\%$ for the sentiment classification task on the IMDB dataset, utilizing the non-linearity and relative positional information in the model. We also achieve $97.0\pm 0.2\%$ accuracy using an astromorphic vision transformer on the CIFAR10 dataset.
IV-A Hardware Specifications & Dataset Details
The IMDB dataset, created by Stanford University, contains $50,000$ movie reviews (each 250-500 words long) labeled as positive or negative. The data is presented in a tab-delimited format, containing fields for movie title, review text, and sentiment label. The CIFAR-10 dataset is a collection of computer vision data, comprising $60,000$ color images with dimensions of $32\times 32$ pixels. These images are categorized into 10 distinct classes, with each class containing 6,000 images.
IV-B Network Architecture
Meanwhile, for image classification, a convolutional and self-attention based neural network (CoAtNet [60]) was utilized to implement our proposed innovations. CoAtNet transformer architecture can best utilize the effectiveness of relative positional information when trained on vision task focused datasets. Furthermore, CoAtNet leverages the capabilities of Convolutional Neural Networks (CNNs) and transformers to enhance its performance. Our architecture implements a $3\times 3$ convolution before creating patches from the images. We use a patch size of $2$ on the image inputs and employ an 8-layered encoder stack with 8-headed self-attention. In addition to the original CoAtNet architecture, we add additional non-linearity and average pooling after encoder layers.
IV-C Performance Evaluation
Table I compares performance of our astromorphic transformer against neuromorphic as well as other non-bioplausible models on the IMDB and CIFAR10 datasets. Our model outperforms RNNs, LSTMs, and GRUs on the IMDB dataset and other bioplausible neuromorphic architectures. However, we did not compare with existing pretrained transformers (BERT, GPT, etc.) because these models are pretrained using non-bioplausible self-attention and have customized pretrained learnable embeddings for natural language processing, whereas we use fixed GloVe embeddings [61] that are simple enough to implement in our astromorphic attention-based transformer architecture. We also test the competitiveness of astromorphic attention in vision transformer architecture. As shown in Table I, our astromorphic vision transformer model competes with artificial and outperforms bioplausible models on the CIFAR10 dataset.
IV-D Ablation Studies
IV-D1 Learning speed
@@
IV-D2 Final accuracy
We implement the theoretical model proposed by [18] to see how it compares without the bioplausibilities proposed herein and find an accuracy of $88.0\pm 0.2\%$ and $87.4\pm 0.1\%$ on the IMDB and CIFAR10 datasets, respectively. This implementation does not incorporate the non-linearity and relative positional information that are integral to our model, which draws inspiration from the dynamics between neurons and astrocytes, thereby underscoring the strong need to consider such bioplausibilities in modeling astromorphic transformers. Our astromorphic transformer also competes with an iso-architecture standard non-bioplausible transformer with softmax attention (see “Transformer” entry in Table 1).
V Conclusion
@@
V-A Appendix A: Transformer Illustration
Transformer employs only encoder architecture for performing sentiment classification and vision tasks. As illustrated in Fig. S5, the input embeddings combined with the positional embeddings enter the encoder layer as input tokens. Then key, query and value are generated from the input tokens and after performing self attention on them, a skip connection is added from the inputs and the overall result is layer normalized and fed into a feed forward network. Another add & normalize operation is performed before feeding them into a linear layer and softmax function to obtain the output probabilities. Sometimes, there might be more than one encoder layer. For different tasks, like machine translation and text generation, encoder-decoder or only decoder type architecture of transformer can be used. However, independent of the type of architecture needed, self-attention is a fundamental mechanism defining transformer operation. The “Multi-Head Attention” Block has self-attention modules parallelly running on different heads obtained from the input embeddings.
Figure S5: Schematic representation of the Transformer Encoder architecture, illustrating the sequential flow of operations. $N$ denotes the number of encoder layers in the network.
V-B Appendix B: Memory Efficient Linearized Attention
Linearized attention involves utilizing feature maps with lower complexity levels, thereby minimizing memory and computational resources. As discussed before, the introduction of non-linearity in biological neurons is facilitated by an activation function ($elu(X)$) in the hidden layer neurons. The utilization of the positive similarity feature map, denoted by $\phi(.)$, characterizes the implementation of the linearized attention operation as described in the work by (Katharopoulos et al. 2020). The utilization of linearized feature attention results in a reduction of computational complexity associated with the computation of the attention score. The advantage of utilizing feature map based linearized attention over the conventional distribution function (i.e. softmax) based self-attention is demonstrated in Fig. S6.
@@ V-B Appendix B: Memory Efficient Linearized Attention
TABLE S2: Hyper-parameter list for the IMDB dataset.
@@
TABLE S3: Hyper-parameter list for the CIFAR10 dataset.
@@
Figure S6: Figure showing the comparison of order of complexity between usual and linearized self-attention. Distribution function can be mapped to softmax function used in usual transformers.
V-C Appendix C: Hyper-parameter Details
The hyper-parameters for the encoder-type astromorphic transformer for the IMDB dataset are listed in Table S2 in the supplementary material. The hyper-parameters for the encoder-type convolution and astromorphic self-attention based transformer for the CIFAR10 dataset are listed in Table S3.